Find the standard deviation of the following data. Round your answer to one decimal place. T01
Answer #1

Solution :

Mean = \mu = \sum X * P(X)

= 0 * 0.1+ 1 *0.1 + 2 *0.2 + 3 * 0.1 + 4 *0.3+5 *0.2

= ( 0+ 0.1+0.4 +0.3+1.2+1 )

= 3

Standard deviation = \sigma

=\sqrt\sumX 2 * P(X) - \mu 2

=  \sqrt[ 02 * 0.1+ 12 *0.1 + 22 *0.2 + 32 * 0.1 + 42 *0.3+52 *0.2 ] -(3)2

= \sqrt [( 0+ 0.1 +0.8 +0.9+4.8+5 ) )] -9

= 11.6 -9

answer=2.6